A coil connected to a 220 V line is used to heat 500.0 g of water at 30°C. It is found out that after 5 minutes, the temperature of the water has risen to 80°C. What is the resistance of the coil? 


1
Expert's answer
2022-04-11T12:02:06-0400

The electrical energy received from the coil equals:


"E=Pt=\frac{U^2}{R}t" ,


where P - electrical power;

U - voltage;

R - resistance;

t - time.


So,


"R=\frac{U^2t}{E}."


Energy, required for heating of the water with mass m from the temperature "T_1" to the temperature "T_2" equals:


"E=cm(T_2-T_1),"


where c - specific heat capacity of the water.


So,


"R=\frac{U^2t}{cm(T_2-T_1)}=\frac{220^2\cdot 300}{4184\cdot 0.5\cdot (80-30)}=138.8\space Ohm."